OpenStudy (anonymous):

If f(x) = -4x - 1, find f(-2).

OpenStudy (kinggeorge):

simply replace x with -2 and solve.

OpenStudy (anonymous):

i keep getting a different answers than what my options are

OpenStudy (kinggeorge):

\[-4(-2) - 1 = 8 - 1 = 7\] so f(-2) = 7

OpenStudy (anonymous):

thank you tons!

OpenStudy (anonymous):

7

OpenStudy (kinggeorge):

Out of curiosity, do you know what you were doing wrong?